OpenStudy (anonymous):

Algebra 2 Word Problem: Set up using 3 variables & 3 equations: "A friend e-mails you the results of a recent high school swim meet. The e-mail states that 24 individuals placed, earning a combined total of 53 points. First place earned 3 points, second place earned 2 points, and third place earned 1 point. There were as many first-place finishers as second and third-place finishers combined." A: Write a system of THREE equations that represents how many people finished in each place. B: How many swimmers finished in first place, second and third?

OpenStudy (ajprincess):

Let the number of swimmers who finished first be x, number of swimmers who finished second be y and number of swimmers who finished third be z. Then x+y+z=24 3x+2y+z=53. Getting this @rosimary?

OpenStudy (anonymous):

Ok Yes - I got that last night while I was working it however I could not figure out equation three at all.

OpenStudy (ajprincess):

There were as many first-place finishers as second & third-place finishers combined." hav u typed the above sentence correctly?

OpenStudy (anonymous):

no I did not hold on let me look at paper

OpenStudy (anonymous):

There were as many first-place finishers as second and third-place finishers combined."

OpenStudy (ajprincess):

What is meant by that sentence is Number of swimmers who finished first=Number of swimmers who finished second+Number of swimmers who finished third x=y+z

OpenStudy (ajprincess):

Nw u hav ur third eqn. Can u solve them to find x, y and z @rosimary?

OpenStudy (anonymous):

give me a moment

OpenStudy (ajprincess):

ha k:)

OpenStudy (anonymous):

I GOT IT!!! x = 12, y = 5 and z = 7

OpenStudy (ajprincess):

yup perfect:)

OpenStudy (anonymous):

Thank you so much. I will try to figure out the rest of my word problems if I get stuck I will open another question. Thanks again.

OpenStudy (ajprincess):

welcome:) and really very good work:)